OpenStudy (anonymous):

the average rate or change of f(x)=5x^2 - 2, from 2 to 4

myininaya (myininaya):

\[\frac{f(4)-f(2)}{4-2}\]

OpenStudy (anonymous):

what do the fs stand for? i get the bottom will be 2

myininaya (myininaya):

Well you have \[f(x)=5x^2-2 \]

myininaya (myininaya):

\[f(4)=5(4)^2-2 \text{ and } f(2)=? \text{ I will let you say what } f(2) \text { is } \]

OpenStudy (anonymous):

5(2)^2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

60/2 ??????????

OpenStudy (anonymous):

or 30/1 simplied. cause they are asking simplified answer.

OpenStudy (anonymous):

or just 30. am i right

myininaya (myininaya):

Yes 60/2 does equal 30